i need to solve each system by graphing and tell wether the system has one solution, infinitely many solutions, or no solution
y=x+1
2x+y=10

Answer by mananth(16946)

-1 x + 1 y = 1 .............1
Total value
2 x + 1 y = 10 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
1 x -1 y = -1
2 x + 1 y = 10
Add the two equations
3 x = 9
/ 3
x = 3
plug value of x in (1)
-1 x + 1 y = 1
-3 + y = 1
y = 1 + 3
y = 4
y = 4
x= 3
y= 4
The solution of system (3,4)
Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when x= 0 y= 1 ,( 0 , 1 )
when y = 0 x= 1 ,( -1 , 0 )
Equation 2
when x= 0 y= 10 ( 0 , 10 )
when y = 0 x= 5 ( 0 , 5 )






m.ananth@hotmail.ca